### v1

# Spatially resolved broad line region in a quasar at z=4

Spatially resolved broad line region in a quasar at z=4
We present the first near-infrared interferometric data of a QSO at z=4. The K-band observations were performed with GRAVITY+ on the VLTI using all 4 UTs, detecting a differential phase signal that traces the spatially resolved kinematics for both the H and H lines in the broad line region. We fit the two lines simultaneously with an updated model that includes distinct rotating and conical outflowing components. We find that more than 80% of the H I line emission from the BLR originates in an outflow with a velocity up to km s-1. This is oriented so that our line of sight is along an edge of the conical structure, which produces the prominent blue wing on the line profile. A combination of anisotropic line emission and mid-plane opacity lead to the single-sided phase signal. The model is able to qualitatively match both the outflowing C IV line profile and the systemic O I fluorescent emission. The derived black hole mass of M⊙ is the highest redshift black hole mass measurement to date obtained directly from BLR dynamics. It is an order of magnitude lower than that inferred from various single epoch scaling relations, and implies that the accretion is highly super-Eddington. With reference to recent simulations, the data suggest that this QSO is emitting close to its radiative limit in a regime where strong outflows are expected around a polar conical region.
Key Words.:
Galaxies: active – quasars: individual: J0529 – quasars: emission lines – quasars: supermassive black holes – Infrared: galaxies1 Introduction
The cosmic epoch at marks a key period between the end of reionisation at (Robertson, 2022; Loeb & Barkana, 2001) and the most active phase of galaxy formation at cosmic noon around (Förster Schreiber & Wuyts, 2020; Tacconi et al., 2020). There is an increasing interest in spatially resolving the molecular and ionised emission lines in galaxies on kiloparsec scales during this epoch (Le Fèvre et al., 2020; Wylezalek et al., 2022; Galbiati et al., 2023; Herrera-Camus et al., 2025; Übler et al., 2024), because doing so can provide insights into the processes that have shaped the universe into what it is now. The role of active galactic nuclei (AGN) is equally important because the black hole accretion rate is rising rapidly during this epoch until it peaks around the same time as galaxy formation (Aird et al., 2015; Kim et al., 2024). Indeed, both the major impact that feedback from AGN can have on galaxy growth (Fabian, 2012; King & Pounds, 2015; Fan et al., 2023), and the existence of local relations between black hole mass and their host galaxy properties (Ferrarese & Merritt, 2000; Gebhardt et al., 2000; Gültekin et al., 2009; McConnell & Ma, 2013; Saglia et al., 2016), indicate that supermassive black holes should co-evolve with galaxies (Kormendy & Ho, 2013; Heckman & Best, 2014).
Observations with JWST, as nicely illustrated by Adamo et al. (2025), have recently highlighted a population of AGN with luminosities overlapping those of local Seyfert nuclei, spanning the entire redshift range from cosmic noon to reionisation (Kocevski et al., 2023; Übler et al., 2023; Greene et al., 2024; Maiolino et al., 2024; Matthee et al., 2024; Suh et al., 2025; Juodžbalis et al., 2025; Scholtz et al., 2025). The data have unexpectedly indicated that, compared to local relations, these black holes are overmassive with respect to their host galaxies. The implication is that they must have grown very rapidly, requiring sustained near-Eddington, or even super-Eddington, accretion, depending on the mass of the seed black holes (Kokorev et al., 2023; Larson et al., 2023; Bogdán et al., 2024; Jeon et al., 2024; Maiolino et al., 2024; Mehta et al., 2024; Suh et al., 2025). However, concerns have been raised about the scale of uncertainties on both the black hole mass and host galaxy mass estimates, and on the use of scaling relations to derive black hole masses (Lambrides et al., 2024; Lupi et al., 2024; Adamo et al., 2025; Bertemes et al., 2025).
Complementary to these efforts are studies focussing on highly luminous QSOs, which are expected to drive the powerful outflows that have the greatest impact on host galaxy evolution (Bischetti et al., 2017). Such QSOs are inferred to have black hole masses M⊙, with a significant fraction close to M⊙, and so are radiating at around their Eddington luminosity (Vietri et al., 2018). The mass outflow rates derived from the prominent broad [O III] lines exceed 1000 M⊙ yr-1, and have kinetic energies around 1% of the bolometric luminosity. Vietri et al. (2018) showed that they follow the expected trend (Sulentic et al., 2011; Sulentic & Marziani, 2015), where those exhibiting prominent [O III] have moderately strong C IV emission with modest blue-shift, while QSOs without such strong [O III] emission have weaker C IV that is more strongly blue-shifted. These authors argued that this could be due to orientation effects affecting observation of the outflows in the broad line region (BLR); and comparison of the C IV blue-shift and FWHM led them to propose that the line profile is a combination of virialised and polar outflowing components. Extremely luminous QSOs at have also been targeted by mid-infrared JWST observations with the aim of resolving the host galaxy (Wylezalek et al., 2022; Vayner et al., 2024), The data have revealed extended [O III] emission in the host galaxies, the outflows, and the circumgalactic medium.
While the properties of outflows on large scales correlate with AGN luminosity, there are also significant departures, for example due to the short timescale variability of AGN which can lead to a wide range of mass loading factors (Zubovas & Nardini, 2020). Similarly, how the outflow is driven, the relative importance of different gas phases, the role of dust, and the degree of clumpiness, all have an impact on the outflow properties and may change with distance from the AGN (King et al., 2011; Faucher-Giguère & Quataert, 2012; Zubovas & King, 2012; Somerville & Davé, 2015; Ward et al., 2024). As such, resolving structures in these massive outflows nearer to their launching sites is a central part of understanding them, but doing so requires additional techniques due to the extremely small scales involved. Velocity resolved reverberation mapping (RM) of nearby AGN has shown that there can be an outflow component to the BLR (Denney et al., 2009; Grier et al., 2013; Du et al., 2016a). More recently, near-infrared interferometry with GRAVITY has been used to spatially resolve the BLR (GRAVITY Collaboration et al., 2018, 2020, 2021), finding that in some local AGN, outflows can dominate the BLR kinematics (GRAVITY Collaboration et al., 2024). The first QSO at has now been observed with the same technique, resolving the BLR kinematics in a highly accreting QSO (Abuter et al., 2024). Observations of several more luminous QSOs at the same epoch have revealed clear signatures of radial flows in a significant fraction of their BLRs (GRAVITY+ Collaboration et al., in prep).
In this paper we take the next step towards higher redshifts, and present interferometric data from GRAVITY+ that resolve the BLR in SMSS J052915.80-435152.0 (hereafter J0529), a highly luminous QSO at z=3.96 (Wolf et al., 2024). Our focus is two-fold. One aspect is to probe the AGN driven outflow on sub-parsec scales, where a broad-line region outflow may be launched. The other is to make a dynamical estimate of the black hole mass, as part of a larger effort to understand whether the scaling relations derived locally can be reliably applied at high redshifts and if needed re-calibrate those relations. Our observations are presented in Sec. 2. We provide a context for our analysis in Sec. 3 by reporting the BLR size and BH mass expected from scaling relations. We then describe our BLR model in Sec. 4 and present our fitting results in Sec. 5, speculating about why they differ so significantly from the scaling relations, and showing how the model can be used to match other emission lines. We summarise our results in Sec. 6.
2 Observations and Data Reduction
The data analysed here were obtained with GRAVITY (GRAVITY Collaboration et al., 2017) as part of programme 114.27UV. They make use of the dual field mode of GRAVITY-wide (GRAVITY+ Collaboration et al., 2022) and the new adaptive optics modules (GRAVITY+ Collaboration, 2025), which are upgrades towards GRAVITY+. As used here, GRAVITY combines the light of all four 8-m UTs of the VLT, to probe spatial scales of milli-arcsec and perform spectro-astrometry at 10 micro-arcsec levels.
The data were obtained on five nights from Sep. to Dec. 2024 with typical seeing (at zenith, and at 500 nm) on different nights of 0.5″ and 0.7″, and coherence timescales of ms. The MEDIUM-COMBINED dual-field wide mode was used, meaning that medium spectral resolution () was used, the polarisations were combined, and the science target and fringe tracking reference star were separate sources with more than 2″ between them. A total of almost 7.8 hr integration was obtained in 64 exposures as shown in Table 1. After processing the data with the GRAVITY instrument pipeline (v. 1.7.0b1), particularly noisy data sets were rejected, reducing the number used from 64 to 59. An additional step was also made to compensate the wavelength dependent variability caused by the dispersion of air in the non-vacuum delay lines as described in Appendix A of GRAVITY Collaboration et al. (2020), using updated components calibrated from GRAVITY-wide observations. Due to the limited signal-to-noise of individual science frames, it was applied on a night-by-night basis: the scaling derived from the combined frames on each night (excluding spectral regions near the H and H lines) was applied to each of those frames individually. Although this approximation cannot take into account changes in delay line length due to tracking, it appears to be sufficient in these circumstances. The resulting differential phase spectra from all nights were then combined with equal weighting, after rejecting (at most a few) deviant values in each spectral channel.
Date | seeing | No. exp. | NDITDIT | |
(″) | (ms) | (s) | ||
27 Sep 2024 | 0.47 | 8.4 | 3 | 4300 |
14 Nov 2024 | 0.72 | 5.9 | 20 | 4100 |
15 Nov 2024 | 0.48 | 10.4 | 22 | 4100 |
12 Dec 2024 | 0.53 | 5.9 | 6 | 4100 |
14 Dec 2024 | 0.72 | 4.7 | 13 | 4100 |
To support the GRAVITY data, an additional K-band spectrum was obtained with the ERIS adaptive optics imager and spectrometer (Davies et al., 2023) as part of programme 112.25M3. ERIS includes an integral field spectrometer (SPIFFIER) which was used here in seeing limited conditions (0.71″ at zenith and 500 nm) with the largest 8″ field of view. The low resolution K-band grating covers 1.96-2.45 m at a resolution of . The data were taken as 11 dithered exposures of 300 s, yielding a total integration time of 55 mins, and reduced with the ERIS instrument pipeline (v 1.6.4). Telluric correction and flux calibration were performed using the B9 IV standard star HD 38343 which has K=7.87 mag. Following this, additional steps were performed to interpolate over the noisiest pixels in the datacube and reduce wavelength dependent variations in the background level. The flux density of the QSO was measured from a spectrum extracted in a large 1.25″ aperture, yielding K = 13.88 mag.
2.1 Image
An image was extracted from the datacube by summing channels in the range 2.028–2.290 m, indicating that the spatial resolution achieved on the QSO itself in the K-band was 0.43″. The image is shown in Fig. 1 on a logarithmic scaling to highlight faint sources. The QSO shows no sub-structure, and the only other source in the field is the M dwarf star previously noted by Wolf et al. (2024) a few arcsec north of the QSO. This strengthens the assessment made by those authors that there is no sign of strong lensing, and hence the luminosity is intrinsic.
2.2 Spectral decomposition
A final spectrum with higher signal-to-noise (S/N) of 85 (but slightly lower flux) than that above was extracted in a 0.62″ aperture better matched to the FWHM. This is used to recover the spectral profile of the broad H I lines, which requires decomposition of the spectrum in order to deblend them from the Fe II features which are known to be strong in highly accreting QSOs (Boroson & Green, 1992; Du et al., 2016b). Since it is already clear from Fig. 2 that the H I lines have a prominent blue wing, we fit these with a pair of Gaussians. This is done solely to match the overall profile and we do not consider the Gaussians individually. We use the same velocity profile for both H I lines, which, although not physically required, is necessary due to limitations of the data. In addition we required zero velocity offset between the H I lines, which was an important constraint on the fit. The reason is that the blue wing of the H line is strongly blended with Fe II and so is primarily constrained by the H line. On the other hand the red side of the H line does provide a primary constraint, whereas the H line is affected by both atmospheric absorption (leading to narrow features and also affecting the slope) and potentially some emission from a blue-shifted wing of [O III] line emission, the peak of which would be just out the spectral range. While we do not have direct evidence for [O III] (and indeed the [O III] line at 4363 Å is not detected although it is expected to be a factor 10-100 weaker), in luminous high-z samples it can be associated with an outflow as indicated by the strong blue-shifted C IV emission described in Sec. 3.2 (Marziani et al., 2017). The intrinsic continuum is expected to be a power law, but over this narrow wavelength range we have adopted a linear function. We consider several templates for the optical Fe II emission, including those based on I Zw 1 (Véron-Cetty et al., 2004), Mrk 493 (Park et al., 2022), a composite based on several QSO spectra (Tsuzuki et al., 2006), and an adaptable template derived from multiple AGN (Kovačević et al., 2010). After trying these options, we adopted the template from Park et al. (2022) because it provides the best match to the Fe II emission on the blue side of the H line; in other regions the templates were equivalently good. Although the properties of the Fe II complex were not tied to the Balmer lines, the broadening of km s-1 FWHM is comparable to the width of the Balmer lines as one would expect if they arise in similar locations (Hu et al., 2015). Finally, specific features in the spectrum that were clearly due to atmospheric effects were excluded during the fitting process. The resulting fit has , and would imply a modest update to the redshift of J0529 as based on the line peak – but see Sec. 5 which indicates the systemic velocity may differ from that. The contribution of the Fe II template and H I line profiles are shown in Fig. 2.
3 Emission lines: inferences and comparisons
From the spectral fit described above, we calculate a line luminosity111Adopting km s-1 Mpc-1, , and ; and using the calculation provided by Wright (2006). of erg s-1, and can extrapolate the optical continuum luminosity to be erg s-1 at 5100 Å. The equivalent width of H is Å, significantly higher than the typical range of Å found for most QSOs (Osterbrock, 1978; Hu et al., 2008).
The line ratio H/H is lower than the value of 0.47 for case B recombination at K (Osterbrock, 1989). It is also marginally lower than the range of values reported by La Mura et al. (2007) for broad line AGN, although it is consistent with the lowest values these authors found for several of their sample. We note that H is within our spectral range, and we find222the S/N of the line ratios is relatively high, in this case , because their shapes are tied to be the same. /H which is slightly less than the lowest values reported by the same authors. As such the line ratios appear physically plausible, and indicate that there might be some modest extinction towards the broad lines.
For the full line profile, we find FWHM km s-1 and km s-1 so that FWHM/. This ratio, has been discussed significantly in the literature (Collin et al., 2006) and is known to vary between AGN. The value we find is significantly lower than the canonical value of 2 often adopted for local AGN when deriving scaling relations (Collin et al., 2006; Woo et al., 2015), and is a direct result of the blue wing on the line profile.
We also calculate the ratio333An alternative that is also commonly used is the ratio of the equivalent widths (Boroson & Green, 1992; Shen & Ho, 2014). In practice it makes little difference because the continuum flux density changes very little between the corresponding wavelengths. where is the flux of Fe II in the range 4434-4684 Å (Du & Wang, 2019). As noted above, this is indicative of the accretion rate and, in a sample of AGN selected to be super-Eddington, Du & Wang (2019) find many of their objects have . Despite the high luminosity of J0529, our value of is modest compared to that sample.
3.1 Single epoch estimates
The estimates above for the optical continuum luminosity and broad line luminosities and widths, enable us to make use of scaling relations to estimate both and . These single epoch estimates provide a context for the dynamical modelling that we perform on the GRAVITY+ data in Sec. 5.
3.1.1 BLR size from scaling relations
Reference | Note | |
1.74 | Bentz et al. (2013) | commonly used |
1.47 | Du et al. (2018) | lower subsample |
1.22 | Du & Wang (2019) | without correction |
0.94 | Du & Wang (2019) | with correction |
0.65 | Woo et al. (2024) | 47 AGN sample |
0.50 | Du et al. (2018) | higher subsample |
0.48 | Woo et al. (2024) | 240 AGN sample |
We have derived the BLR size from 4 different relations and summarised them in Table 2. One of the best known relations between the AGN luminosity and size of the BLR is that of Bentz et al. (2013). The AGN luminosities were derived using a surface brightness decomposition to correct for the host galaxy; and the BLR sizes were derived from a simple measure of the time lag obtained from RM. An emphasis to AGN with high accretion rates was given by Du et al. (2018), who derived different relations depending on the accretion rate. We show here the relation for both lower and higher accretion rates, which yield very different results. We also consider a version of that relation which was updated by Du & Wang (2019) to include an explicit correction for the accretion rate via the parameter. For J0529 this provides a substantial reduction to the inferred BLR size compared to Bentz et al. (2013), although for J0529 the correction itself is only a 25% effect. Finally, we consider estimates from Woo et al. (2024) who provided two fits. We include both, one based on a combination of 47 AGN from two specific samples, and one derived from 240 AGN that are mostly from different sources in the literature. These authors found no strong dependent deviation from a relation, but they did note that the slope of the relation is likely to be shallower than that popularly used. These differences highlight the uncertainties associated with employing scaling relations. Since there is also a scatter (typically reported as dex) associated with each relation, here we can conclude only that is likely in the range 0.5-1.7 pc.
Reference | Note | |
10.92 | Dalla Bontà et al. (2020) | Eq. 38 using & LHβ |
10.24 | Wolf et al. (2024) | continuum modelling; also C IV & Mg II |
10.16 | Woo et al. (2015) | Eq. 2 using & L5100 |
9.87 | Reines et al. (2013) | Eqs. 2, 3, & 5; converting FWHMHβ & L5100 to H |
9.58 | Vietri et al. (2020); Bongiorno et al. (2014) | Eq. 4 or Eq. 2 respectively, using FWHMHβ & L5100 |
3.1.2 Black hole mass from scaling relations
In a similar approach, we estimate the black hole mass from several different relations which are summarised in Table 3. The mass reported by Wolf et al. (2024) in their discovery paper was based on both modelling of the continuum shape and use of the C IV 1549 and Mg II 2799 line widths. A recent re-assessment of the use of H to estimate was provided in Eq. 38 of Dalla Bontà et al. (2020), who included a (statistical) correction for Eddington ratio . These authors also argued that the dispersion is a less biassed tracer of the line width than FWHM (although use of for profiles with Lorentzian-like wings also poses difficulties, Santos et al. 2025); and prefer to use the line rather than continuum luminosity in order to avoid bias from core dominated radio sources. An important part of the relation is converting the virial mass into a black hole mass via the parameter, and Dalla Bontà et al. (2020) adopt from Batiste et al. (2017). For J0529, the inferred BH mass is . A likely reason for such a high mass is the bias on both the line dispersion and luminosity caused by the blue wing. Indeed, using only the Gaussian corresponding to the main part of the line profile yields a lower mass consistent with that of Wolf et al. (2024). An earlier relation which did not attempt to make any corrections, and was based on the relation of Bentz et al. (2013), is given as Eq. 2 of Woo et al. (2015), for which they adopt (a very different value than used by Dalla Bontà et al. (2020), emphasizing that it depends on how the scaling relation is set up and whether FWHM or is used). From this relation we infer a value of using the full line profile. Reines et al. (2013) derived a relation based on the continuum luminosity and FWHM of the H line, providing conversions for these quantities to H luminosity and FWHM in Eqs. 2 and 3. This also builds on the relation of Bentz et al. (2013), and uses the conversions to H from Greene & Ho (2005). We can put our measured values into those equations, apply the conversions, and hence estimate a mass from their Eq. 5. Doing this, and noting that they use , yields . The final relation we consider is Eq. 4 from Vietri et al. (2020) which was applied to luminous QSOs at , and which was taken directly from Eq. 2 in Bongiorno et al. (2014) where it was applied to X-ray obscured red QSOs at . As for the previous relation it uses FWHMHβ and L5100, yielding . The conclusion here is that is associated with even more uncertainty than , and that there may be a systematic difference when using vs FWHM. For J0529 the scaling relations tend to give values around .
3.2 Comparison to UV lines
In this section we compare the H I line profile to the UV lines C IV and O I in the Xshooter spectrum published by Wolf et al. (2024). To do so, we have subtracted an estimate of the true continuum level as indicated by the lowest regions in the spectral ranges around the two lines. We have not tried to fit the iron emission, nor other neighbouring emission, since the aim is only to compare the line profiles.
The right panel of Fig. 3 shows the comparison to O I 1302 Å, a low excitation line that in AGN is expected to arise in the BLR through Ly fluorescence (Netzer & Davidson, 1979; Kwan & Krolik, 1981; Matsuoka et al., 2007). O I 8446 Å is also enhanced through this process and the ratio between these lines in the resulting radiative cascade can provide a measure of the extinction towards the BLR (Netzer & Davidson, 1979). Given that the Balmer line ratios in Sec. 3 hint that there might be some modest extinction, an independent measure like this would enable one to clarify the impact. There is a good match on the red side of the profiles, but the O I line shows no evidence for a blue wing suggesting it is more likely associated with the systemic velocity of the QSO. We note that the peak of the O I line is contaminated by deep absorption across a range of velocities, the cause of which is unclear.
The C IV 1549 Å line in the left panel presents a contrasting picture since it is only associated with the outflow, and can be traced to a velocity of . It confirms the nuclear nature of the outflow, and from the size-luminosity relation of the C IV line and 1350 Å continuum (based on 7-13 yr RM campaigns) we infer a characteristic radius for the C IV emission of 0.4 pc (Kaspi et al., 2021; Shen et al., 2024). Analysis of the C IV line is not our main aim, and it is often done in association with [O III] emission which is outside our spectral range. We therefore note only that the line properties are often grouped into two classes of QSOs related to whether the C IV is dominated by a more systemic or strongly blue-shifted component, which are also broadly linked to the properties of Narrow Line Seyfert 1 galaxies (Sulentic et al., 2011; Sulentic & Marziani, 2015; Marziani et al., 2017; Vietri et al., 2018, 2020). In this scheme the moderate equivalent width, blue-shifted C IV line in J0529 would be associated with a disk wind.
Because the C IV line is easily observable in high redshift QSOs and is known to originate in the BLR, it has been used to estimate despite the often unusual line properties (Vestergaard & Peterson, 2006; Park et al., 2013; Runnoe et al., 2013). While this is physically reasonable for QSOs in which the line is both approximately symmetric and close to the systemic velocity, it is harder to explain a relation between the line properties and for emission originating in an outflow. Despite this, Coatman et al. (2017) found an empirical relation between the line’s blue-shift and a correction factor that can be applied to the inferred mass, to match it to other scaling relations. However, Temple et al. (2023) have shown that the situation is more complex, and that there is a 2-dimensional parameter space relating the blue-shift (and equivalent width) to a combination of and . They showed there is a wedge in that parameter space where the C IV blue-shift increases (and equivalent width decreases), and that in general the highest blue-shifts are observed only in AGN with both and . Their modelling to reproduce these effects is bringing us to a better understanding of the cause of the two populations of QSOs with differing line properties noted above. J0529 has a blue-shift km -1 higher than their sample range, and has both a high and , consistent with their analysis.
4 Description of the BLR model
We make use of the phenomenological model described by Pancoast et al. (2014), our implementation of which is given in detail in GRAVITY Collaboration et al. (2020) with summaries in GRAVITY Collaboration et al. (2021) and GRAVITY Collaboration et al. (2024). The model implements the BLR as an ensemble of clouds moving in the potential of the central black hole. Physically, clouds in the BLR are likely to be transient objects, plausibly formed by shock compression due to gas collisions on small scales (as has been adopted to model the ISM on scales of -10 pc in AGN, Vollmer et al. 2018). However, irrespective of the true nature of the clouds, in our phenomenological model the use of the term is for convenience. The clouds should be taken to represent line emitting test particles rather than physical entities, because the model parameterizes the distribution and kinematics of the line emission rather than the entire gas content of the BLR.
The model of Pancoast et al. (2014) is a good representation of the physics motivating the FRADO (Failed Radiatively Accelerated Disk Outflow) model, proposed by Czerny & Hryniewicz (2011) and developed by Czerny et al. (2017) and Baskin & Laor (2018). In this concept, the anisotropic emission from the accretion disk and the effects of shielding enable dust to exist in a thin disk at radii smaller than the nominal sublimation radius. The radiation pressure from the hot disk can accelerate a dusty gas cloud vertically until it has a direct line of sight to the accretion disk. At this point the dust sublimates, and the cloud continues ballistically before falling back down to the disk. The BLR therefore forms a flared region above the disk plane, as proposed by other authors (Goad et al., 2012; Ramolla et al., 2018; Hopkins, 2025). These models are consistent with the dust having a bowl-shaped form that was proposed to explain the difference in dust sizes between interferometric and reverberation mapping measurements (Gravity Collaboration et al., 2024).
A limitation of the FRADO model is that it only considers vertical motion of the clumps, and does not address their radial motion. Once clouds are driven high enough that the dust sublimates, the direct line of the sight they have to the AGN is also expected to lead to an outward acceleration. This may be either by UV-line driving, as is believed to be happening for ultra fast outflows (UFOs) from the accretion disk (Mizumoto et al., 2021), or as a magneto-centrifugal wind (Emmering et al., 1992), which has also been proposed as an outflow mechanism on 1-10 pc scales (Vollmer et al., 2018). The BLR outflow will form part of a hierarchy of outflows, located between the small scale UFOs and the radiatively driven outflow that accelerates dusty gas to create geometrically thick obscuration (Ramos Almeida & Ricci, 2017; Chan & Krolik, 2017; Hönig, 2019). Physically, and based on simulations (Kudoh et al., 2023, 2024; Hopkins et al., 2025; Hopkins, 2025), one would expect outflow to be driven at the surface of the BLR and inflow to occur closer to the disk plane. However, in the Pancoast et al. (2014) model, radial motions are included in a way that allows a fraction of the clouds throughout the bulk of the BLR to be either all inflowing or all outflowing (and unbound) on trajectories that are essentially radial.
In order to better physically motivate the way radial flows are handled in our phenomenological model, we have updated its implementation for the analysis here, while keeping the model as simple as possible. The original flared disk model is used to represent the static part of the BLR described by Czerny et al. (2017) and Baskin & Laor (2018). It is unchanged except that outflows are excluded, while inflow is in principle allowed. However, after initial iterations of the model, it was clear that for J0529 the fraction of inflowing clouds was small enough that, for simplicity, it could be excluded entirely with negligible impact. Thus, in this case, the motions in the disky part of the model are purely rotational.
A specific outflow component is added as an additional wedge that extends above the upper surface of that flared disk as shown in Fig. 4. Again for simplicity, to avoid including more free parameters, we assume that the outflow and the disk components are spatially connected. It represents the concept put forward by Elvis (2000) as well as the outflowing part in the simulations of Kudoh et al. (2023, 2024) and Hopkins et al. (2025), and has been modelled explicitly by Yong et al. (2017, 2020). The total angular height of the BLR is then the sum of for the flared disk and for the outflow. Thus a situation where would mean that there are clouds on orbits perpendicular to the disk plane.
Clouds in the outflow are required to be rotating around the central black hole with the same circular velocity they would have if they were in the midplane at the same projected radius. They must also be outflowing with a radial velocity . The original model defined the outflow velocity at a radius to be the same as the escape velocity at that radius. A couple of examples show that this was a reasonable assumption. For and derived in NGC 3783 (GRAVITY Collaboration et al., 2021), where about half the clouds are on unbound orbits, km s-1 at the average radius of the BLR is only a factor greater more than the 1800 km s-1 FWHM and within the full line profile; Similarly, for the nominal size (1 pc) and mass ( M⊙) inferred in Sec. 3.1 for J0529, V km s-1 is only a factor of a few greater than the dispersion of the line profile including the blue wing. However, the assumption that is also a limitation, and can lead to unphysical geometries for prominent fast outflows for which the model is not suited. Rather than follow the same prescription for we apply an improvement suggested by Pancoast et al. (2014) to link it to the escape velocity at the wind launching radius (Murray et al., 1995; Long et al., 2023). Our implementation ties Vout more flexibly to the escape velocity at the minimum radius and still enables it to decrease with radius:
where is the minimum radius for cloud distribution, is a scaling factor (expected to be in the range one to a few) for the escape velocity at , and is a negative index in the range allowing to decrease with increasing radius. We argue that such a radial decrease should be expected because reverberation mapping of the C IV line, which has much a higher excitation potential than H I, tends to yield a smaller radial scale than H (Lira et al., 2018; Shen et al., 2024); and the line is often blue-shifted with respect to the H I lines (Coatman et al., 2017; Temple et al., 2023), as also illustrated for J0529 in Fig. 3. Thus it is reasonable to conclude that the outflow velocity is higher at smaller radii.
Finally, and again to keep the number of free parameters as low as possible, the vertical angular distribution of clouds in each of the outflowing and rotating regions is fixed to be uniform (rather than, as done in the original model, using to control whether the clouds are concentrated more towards the top edge of the distribution). However, because there is a new parameter defining the fraction of clouds in the outflow, it is possible for the cloud density to differ between the outflowing versus rotating regions.
5 The fitted BLR model
We fit the H and H lines simultaneously in a similar way to that described in Kuhn et al. (2024): all the parameters in the fit are tied to be the same for both lines except for the two that define their radial distribution. These are the parameter that defines the form of the distribution, and the scale factor which is linked to the average radius such that , where is the same minimum radius for both lines. The details of the fit are given in Appendix A, including Table 4 summarising the best fitting parameters corresponding to the model, and Fig. 8 of the resulting spectra and differential phases for each baseline. As a complement to this, the top two rows of Fig. 5 show the data and fitted model for each of the line profiles as well as the sum of the phases from baselines UT 4-2, 4-1, and 3-2 which contain the strongest signal. The three main points from the phase data in Fig. 5 are (i) the phase signal for H is about 1/3 the strength of that for the H line, as expected from their equivalent widths; (ii) the phase signal is single-sided which is indicative of predominant outflow rather than the S-shape associated with rotation; and (iii) there are two distinct peaks in the phase signal, which are related to the separate outflowing and disky components. This last point is developed further in the bottom panels of Fig. 5 which show the contribution of the outflowing and disky components to the total line profile. It is clear that emission from the outflow is not just limited to the blue wing of the profile, but makes up a substantial part of the core as well. In contrast the disky component has only a subdominant contribution on the red side of the line. As such, the model implies that the systemic velocity of J0529, which should be traced by the disky component of the BLR that peaks at 2.423 m, is offset by about 1200 km s-1 from the peak of the line. This would indicate an update of the redshift to .
Histograms of the posterior distributions for all the parameters are presented in Fig. 9, and the most relevant results for our analysis are additionally shown in the corner plot in Fig. 10. These figures indicate that the fit is well constrained; and that while other potential solutions might exist, these are at a much lower probability.
5.1 BLR geometry and size
The fitted parameters summarised in Table 4 show that % of the clouds are in the outflow, and that this component therefore completely dominates the BLR. The remaining 17% of the clouds are in the disky component on bound elliptical orbits. Our line of sight is near the top edge of the outflow, because the inclination is from face-on while the total angle from the mid-plane subtended by BLR clouds is . We note that if the line of sight was oriented further into the outflowing wedge, one might expect to see broad absorption lines in the rest-frame optical spectrum (Elvis, 2000) but these are not observed (Wolf et al., 2024). The emission is anisotropic with indicating that the side of the clouds illuminated by the AGN emits preferentially. Equally important is the mid-plane opacity with meaning that the emission from clouds behind the mid-plane is almost entirely blocked, consistent both with the basis of the FRADO model in which a dusty disk extends to small radii at low scale heights and also the optically thick nature of super-Eddington accretion disks. Together these lead to the configuration in Fig. 6, where the clouds are coloured according to their line-of-sight velocity and have a size corresponding to their apparent brightness.
The average radius for the BLR is pc for H and pc for H. While these are the same within their uncertainties (see Table 4 and Fig. 9), this difference is expected both observationally, based on RM delays for different lines (Bentz et al., 2010), and on modelling of radiation pressure confined BLR clouds (Netzer, 2020).
While the size appears to match that inferred by applying a correction to the R-L relation favoured by Du & Wang (2019), this does not provide an explanation of why is so low. Indeed, from the plot of FWHM versus in Shen & Ho (2014), both of these values for J0529 seem rather typical of QSOs generally. Instead we propose that much of the deviation from the scaling relations that has previously been attributed to luminosity or accretion rate, could, analogous to the C IV line, be due to the impact of the outflow on the H I line properties. Indeed, the disky part of the BLR in Fig. 5 contains only 27% of the total line luminosity. Applying this to the typical equation which has a typical form where is usually taken to represent the total AGN luminosity. If instead we consider only the luminosity associated with the disky part of the BLR tracing rotationally supported gas (using the line luminosity as a suitable proxy), then the inferred size is reduced by a factor . This would bring the estimate from Bentz et al. (2013) into good agreement with our measurement.
Similarly, using only the disky part of the BLR would bring the equivalent width down to Å, within the range of typical values. And we would find which is what one would expect for a highly accreting QSO, particularly in the context of Eddington ratio discussed below. However, applying this new value would lead to a revised BLR size from Du & Wang (2019) that is much too small. We therefore speculate that the over-estimate of the BLR size that has been discussed in the literature for highly accreting and highly luminous QSOs could be related to the impact of BLR outflows on the line properties.
5.2 Dynamical black hole mass
The dynamical black hole mass we derive is M⊙, which is lower than those inferred from scaling relations by an order of magnitude. The low mass is largely driven by the small rotational velocity required in the model, which has to be less than the line width in order to allow for the outflow.
The form of the scaling relation used to estimate black hole mass is typically where represents AGN luminosity and (or FWHM) the line width. If, as above, we take to be a proxy for the line luminosity (noting that line luminosity is used explicitly in some relations), then applying this to only the disky part of the BLR profile which has km -1 and contains only 27% of the luminosity, would reduce the inferred by a factor 14 (or a factor 5 if using FWHM, due to the differing FWHM/ ratios). This use of the rotational velocity of the BLR rather than the full line width (which is broadened by the outflow) would bring all the single epoch estimates much closer to the BH mass we have measured, and supports the speculation above that the outflow could be a major contributor to the bias in single epoch estimates for luminous QSOs
5.3 Eddington ratio
The Eddington ratio is determined by the black-hole mass and the bolometric luminosity , which is commonly derived from a monochromatic luminosity and a bolometric correction . Also, the emission from the accretion disc requires an anisotropy correction that depends on the viewing angle, and to a smaller extent on the black-hole spin since the deflection of light due to gravitational lensing is stronger for disc annuli that are closer to the black hole (Nemmen & Brotherton, 2010; Lai et al., 2023). Generic values based on a single average spectral energy distribution (SED; Richards et al. 2006) are usually appropriate when the disc is large compared to the innermost stable circular orbit (ISCO) around the black hole. While the overall size and surface area of a disc is set by its luminosity, black holes of increasing mass will have the hole inside the ISCO grow proportionally to mass. Discs around higher-mass black holes thus lack the hottest and most UV-emitting parts, which are otherwise present in discs of similar luminosity around lower-mass black holes. This implies a non-linear dependence of on mass and disc luminosity, which was explored in the framework of thin-disc models by several authors (e.g. Nemmen & Brotherton, 2010; Netzer, 2019; Lai et al., 2023). These explorations tend to offer a mean relation for as a function of monochromatic luminosity assuming a mean related black-hole mass or Eddington ratio.
In our case of J0529, we use the 5100 Å luminosity from Sec. 3 and the 33∘ inclination implied by our model. The mean relations by the three aforementioned works result in values ranging from 2.2 to 2.7, but are not applicable in the case of J0529. Such low values correspond to discs around black holes with masses of , whose SEDs visibly turn down towards the far-UV part of the spectrum. For example, Lai et al. (2023) show the red UV spectrum of their object J2157 that is in stark contrast to the blue spectrum of J0529, which implies a larger value for J0529 and suggests a lower-mass black hole for common stationary disc models. Also, using the mean relation from Netzer (2019), the luminosity of J0529 puts it far outside of the parameter space covered by the physically allowed range of black hole spin, while extrapolating their model calculation for any allowed spin to the correct luminosity puts it not only into a clearly super-Eddington regime but also suggests .
[…]
5.4 Predictions of UV line profiles
Using our BLR model we can predict the spectral line profile of the O I and C IV lines that were presented in Sec. 3.2. We have done this in a simplified way because our model is based on dynamics rather than photoionisation, hence we infer the location at which the corresponding line emission is produced by adjusting the size of the BLR for each line based on the ratio of the time lag of that line with respect to H as measured empirically from RM campaigns. For O I the scaling is because, as noted in Sec. 3.2, it is produced by Ly fluorescence and so is expected to arise at a similar location to the H I lines. Other low excitation lines such as Mg II and Fe II also have broadly similar time lags to H (Maoz et al., 1993; Hu et al., 2015; Prince et al., 2023; Shen et al., 2024). For simplicity we also assume that the O I line originates in the disky part of the BLR. For C IV, we adopt a time lag that is 30% of that for H, although there is again a large range of variation and the ratio may depend on luminosity (Lira et al., 2018; Shen et al., 2024). In addition the shape of the cross-correlation function may differ significantly between the lines (Horne et al., 2021). We assume C IV arises only in the outflowing part of our model. The resulting line profiles in Fig. 7 are similar to what one might expect given the contributions of disky and outflowing regions to the H line in Fig. 5. However, an important difference is the stronger and more extended blue wing on the predicted C IV profile, which qualitatively matches the observation shown in the lower panel of Fig. 3, and is a result of the way that decreases with radius. The higher excitation C IV originates from smaller radii and, lending support to the decreasing velocity profile we have adopted, has a higher outflow velocity.
5.5 Outflow rate
Following Shimizu et al. (2019), we estimate the outflow rate from our model as
where is the radial range over which the estimate is made. We have applied this in a simplified way to calculate a single outflow rate between and the mean radius at 0.9 pc. For we use the mean radial velocity of the clouds within this range, and we assume a canonical density of cm-3. The resulting outflow rate, based on the inferred H luminosity, is M⊙ yr-1. We can apply a factor 2 correction to compensate the anisotropy in the model and another factor 2 because of the mid-plane opacity to account for the reverse outflow, increasing the implied outflow rate to M⊙ yr-1. This is very small compared to the accretion rate of M⊙ yr-1 required to maintain the luminosity, assuming a radiative efficiency of for super-Eddington sources (Jiang et al., 2019). However, it represents only the outflow of ionised gas in the BLR, and does not include the accretion disk wind launched on smaller scales, nor the radiation driven dusty wind from larger scales (which has the additional impact of reducing the inflowing mass that reaches the accretion disk). The lower outflow rate associated with the BLR is consistent with what was found by Vietri et al. (2018) for highly luminous QSOs. These authors showed that the kinetic power of the outflow from C IV in the BLR was an order of magnitude or more lower than that from [O III] on kiloparsec scales; and that winds from the accretion disk dominate the outflow on small scales to make up the difference.
6 Conclusions
We have used the VLTI interferometer GRAVITY+ to spatially resolve the kinematics of the BLR of the luminous z=4 QSO SMSS J052915.80-435152.0. We measured the differential phase spectra across both the H and H lines and combined these with a detailed flux spectrum from ERIS. We applied an updated modelling procedure that has a disky component and a separate outflowing component; and we fit both the flux profiles and phase spectra of the two H I lines simultaneously, allowing only the radial profiles of their line emission distribution to vary. The systemic velocity in the model differs by km s-1 from peak of the broad line emission, implying a redshift for the QSO of . Based on the model, we conclude:
The BLR is resolved to be about 1 pc in radius and is dominated by outflow, with velocities up to km s-1 which create the prominent blue wing and also contribute to much of the line profile’s core. Both high mid-plane opacity and anisotropic emission lead to the observed single-sided differential phase profile. Our model can plausibly predict both the observed blue-shifted C IV 1549 Å profile and the systemic O I 1302 Å profile.
The black hole mass of M⊙ is the highest redshift measurement obtained directly from spatially resolved BLR kinematics to date. It is an order of magnitude lower than the masses inferred from a variety of scaling relations.
The QSO has a luminosity of , which recent simulations indicate is plausible for accretion rates in a range around – close to, but not yet, saturating the luminosity; and in a regime where strong polar conical outflows are expected.
The ratio (of the strength of the iron to H emission) would imply a rather modest accretion rate, contrary to other evidence indicating there is a high accretion rate. We speculate that this discrepancy might be attributable to the dominance of the outflow in the BLR. If the scaling relations are applied using only the disky part of the BLR profile, the BH mass inferred from scaling relations is reduced, and the implied accretion rate increases. However, by comparing Fig. 2 and the bottom panels of Fig. 5 we caution that the decomposition of the BLR profile into outflowing and disky parts is dependent on geometry and anisotropy parameters, so cannot be done in a naive way by fitting a pair of Gaussian profiles.
An estimate of the outflow rate of ionised gas in the BLR is M⊙ yr-1. This is only a few percent of the accretion rate, indicating it is only a minor part of the total outflow rate which includes winds launched on both smaller and larger scales.
Acknowledgements.
We thank the staff of Paranal Observatory for their enthusiasm and help during the visitor runs in which the observations here were performed. This paper has benefitted from useful and lively discussion with many people on the BLR and outflows. In particular RD thanks B. Vollmer and K. Tristram, as well as the participants of the 2025 Vasto Accretion Meeting; RD and TS thank H. Netzer; JS thanks K. Inayoshi, J. Dai, and C. Jin for insightful suggestions on the BLR outflow modeling. JS is supported by The Fundamental Research Funds for the Central Universities, Peking University (7100604896) and the China Manned Space Program with grant no. CMS-CSST-2025-A09. LCH was supported by the National Science Foundation of China (12233001) and the China Manned Space Program (CMS-CSST-2025-A09). J.S-B. acknowledges the support received by the UNAM DGAPA-PAPIIT project AG-101025 and to the SECIHTI Ciencia de Frontera project CBF-2025-I-3033.References
- Abramowicz et al. (1988) Abramowicz, M. A., Czerny, B., Lasota, J. P., & Szuszkiewicz, E. 1988, ApJ, 332, 646
- Abuter et al. (2024) Abuter, R., Allouche, F., Amorim, A., et al. 2024, Nature, 627, 281
- Adamo et al. (2025) Adamo, A., Atek, H., Bagley, M. B., et al. 2025, Nature Astronomy [arXiv:2405.21054]
- Aird et al. (2015) Aird, J., Coil, A. L., Georgakakis, A., et al. 2015, MNRAS, 451, 1892
- Baskin & Laor (2018) Baskin, A. & Laor, A. 2018, MNRAS, 474, 1970
- Batiste et al. (2017) Batiste, M., Bentz, M. C., Raimundo, S. I., Vestergaard, M., & Onken, C. A. 2017, ApJ, 838, L10
- Bentz et al. (2013) Bentz, M. C., Denney, K. D., Grier, C. J., et al. 2013, ApJ, 767, 149
- Bentz et al. (2010) Bentz, M. C., Walsh, J. L., Barth, A. J., et al. 2010, ApJ, 716, 993
- Bertemes et al. (2025) Bertemes, C., Wylezalek, D., Rupke, D. S. N., et al. 2025, A&A, 693, A176
- Bischetti et al. (2017) Bischetti, M., Piconcelli, E., Vietri, G., et al. 2017, A&A, 598, A122
- Bogdán et al. (2024) Bogdán, Á., Goulding, A. D., Natarajan, P., et al. 2024, Nature Astronomy, 8, 126
- Bongiorno et al. (2014) Bongiorno, A., Maiolino, R., Brusa, M., et al. 2014, MNRAS, 443, 2077
- Boroson & Green (1992) Boroson, T. A. & Green, R. F. 1992, ApJS, 80, 109
- Chan & Krolik (2017) Chan, C.-H. & Krolik, J. H. 2017, ApJ, 843, 58
- Coatman et al. (2017) Coatman, L., Hewett, P. C., Banerji, M., et al. 2017, MNRAS, 465, 2120
- Collin et al. (2006) Collin, S., Kawaguchi, T., Peterson, B. M., & Vestergaard, M. 2006, A&A, 456, 75
- Czerny & Hryniewicz (2011) Czerny, B. & Hryniewicz, K. 2011, A&A, 525, L8
- Czerny et al. (2017) Czerny, B., Li, Y.-R., Hryniewicz, K., et al. 2017, ApJ, 846, 154
- Dalla Bontà et al. (2020) Dalla Bontà, E., Peterson, B. M., Bentz, M. C., et al. 2020, ApJ, 903, 112
- Davies et al. (2023) Davies, R., Absil, O., Agapito, G., et al. 2023, A&A, 674, A207
- Denney et al. (2009) Denney, K. D., Peterson, B. M., Pogge, R. W., et al. 2009, ApJ, 704, L80
- Du et al. (2016a) Du, P., Lu, K.-X., Hu, C., et al. 2016a, ApJ, 820, 27
- Du et al. (2016b) Du, P., Lu, K.-X., Zhang, Z.-X., et al. 2016b, ApJ, 825, 126
- Du & Wang (2019) Du, P. & Wang, J.-M. 2019, ApJ, 886, 42
- Du et al. (2018) Du, P., Zhang, Z.-X., Wang, K., et al. 2018, ApJ, 856, 6
- Elvis (2000) Elvis, M. 2000, ApJ, 545, 63
- Emmering et al. (1992) Emmering, R. T., Blandford, R. D., & Shlosman, I. 1992, ApJ, 385, 460
- Fabian (2012) Fabian, A. C. 2012, ARA&A, 50, 455
- Fan et al. (2023) Fan, X., Bañados, E., & Simcoe, R. A. 2023, ARA&A, 61, 373
- Faucher-Giguère & Quataert (2012) Faucher-Giguère, C.-A. & Quataert, E. 2012, MNRAS, 425, 605
- Ferrarese & Merritt (2000) Ferrarese, L. & Merritt, D. 2000, ApJ, 539, L9
- Förster Schreiber & Wuyts (2020) Förster Schreiber, N. M. & Wuyts, S. 2020, ARA&A, 58, 661
- Galbiati et al. (2023) Galbiati, M., Fumagalli, M., Fossati, M., et al. 2023, MNRAS, 524, 3474
- Gebhardt et al. (2000) Gebhardt, K., Bender, R., Bower, G., et al. 2000, ApJ, 539, L13
- Goad et al. (2012) Goad, M. R., Korista, K. T., & Ruff, A. J. 2012, MNRAS, 426, 3086
- GRAVITY+ Collaboration (2025) GRAVITY+ Collaboration. 2025, A&A, submitted
- GRAVITY Collaboration et al. (2017) GRAVITY Collaboration, Abuter, R., Accardo, M., et al. 2017, A&A, 602, A94
- GRAVITY+ Collaboration et al. (2022) GRAVITY+ Collaboration, Abuter, R., Allouche, F., et al. 2022, A&A, 665, A75
- GRAVITY Collaboration et al. (2021) GRAVITY Collaboration, Amorim, A., Bauböck, M., et al. 2021, A&A, 648, A117
- GRAVITY Collaboration et al. (2020) GRAVITY Collaboration, Amorim, A., Bauböck, M., et al. 2020, A&A, 643, A154
- GRAVITY Collaboration et al. (2024) GRAVITY Collaboration, Amorim, A., Bourdarot, G., et al. 2024, A&A, 684, A167
- Gravity Collaboration et al. (2024) Gravity Collaboration, Amorim, A., Bourdarot, G., et al. 2024, A&A, 690, A76
- GRAVITY Collaboration et al. (2018) GRAVITY Collaboration, Sturm, E., Dexter, J., et al. 2018, Nature, 563, 657
- Greene & Ho (2005) Greene, J. E. & Ho, L. C. 2005, ApJ, 630, 122
- Greene et al. (2024) Greene, J. E., Labbe, I., Goulding, A. D., et al. 2024, ApJ, 964, 39
- Grier et al. (2013) Grier, C. J., Peterson, B. M., Horne, K., et al. 2013, ApJ, 764, 47
- Gültekin et al. (2009) Gültekin, K., Richstone, D. O., Gebhardt, K., et al. 2009, ApJ, 698, 198
- Heckman & Best (2014) Heckman, T. M. & Best, P. N. 2014, ARA&A, 52, 589
- Herrera-Camus et al. (2025) Herrera-Camus, R., González-López, J., Förster Schreiber, N., et al. 2025, A&A, 699, A80
- Hönig (2019) Hönig, S. F. 2019, ApJ, 884, 171
- Hopkins (2025) Hopkins, P. F. 2025, The Open Journal of Astrophysics, 8, 56
- Hopkins et al. (2025) Hopkins, P. F., Su, K.-Y., Murray, N., et al. 2025, The Open Journal of Astrophysics, 8, 48
- Horne et al. (2021) Horne, K., De Rosa, G., Peterson, B. M., et al. 2021, ApJ, 907, 76
- Hu et al. (2015) Hu, C., Du, P., Lu, K.-X., et al. 2015, ApJ, 804, 138
- Hu et al. (2008) Hu, C., Wang, J.-M., Ho, L. C., et al. 2008, ApJ, 683, L115
- Inayoshi et al. (2020) Inayoshi, K., Visbal, E., & Haiman, Z. 2020, ARA&A, 58, 27
- Jeon et al. (2024) Jeon, J., Bromm, V., Liu, B., & Finkelstein, S. 2024, in 45th COSPAR Scientific Assembly, Vol. 45, 1415
- Jiang & Dai (2024) Jiang, Y.-F. & Dai, L. 2024, arXiv e-prints, arXiv:2408.16856
- Jiang et al. (2019) Jiang, Y.-F., Stone, J. M., & Davis, S. W. 2019, ApJ, 880, 67
- Juodžbalis et al. (2025) Juodžbalis, I., Maiolino, R., Baker, W. M., et al. 2025, arXiv e-prints, arXiv:2504.03551
- Kaspi et al. (2021) Kaspi, S., Brandt, W. N., Maoz, D., et al. 2021, ApJ, 915, 129
- Kim et al. (2024) Kim, S. J., Goto, T., Ling, C.-T., et al. 2024, MNRAS, 527, 5525
- King & Pounds (2015) King, A. & Pounds, K. 2015, ARA&A, 53, 115
- King et al. (2011) King, A. R., Zubovas, K., & Power, C. 2011, MNRAS, 415, L6
- Kocevski et al. (2023) Kocevski, D. D., Onoue, M., Inayoshi, K., et al. 2023, ApJ, 954, L4
- Kokorev et al. (2023) Kokorev, V., Fujimoto, S., Labbe, I., et al. 2023, ApJ, 957, L7
- Kormendy & Ho (2013) Kormendy, J. & Ho, L. C. 2013, ARA&A, 51, 511
- Kovačević et al. (2010) Kovačević, J., Popović, L. Č., & Dimitrijević, M. S. 2010, ApJS, 189, 15
- Kudoh et al. (2023) Kudoh, Y., Wada, K., Kawakatu, N., & Nomura, M. 2023, ApJ, 950, 72
- Kudoh et al. (2024) Kudoh, Y., Wada, K., Kawakatu, N., & Nomura, M. 2024, ApJ, 977, 48
- Kuhn et al. (2024) Kuhn, L., Shangguan, J., Davies, R., et al. 2024, A&A, 684, A52
- Kwan & Krolik (1981) Kwan, J. & Krolik, J. H. 1981, ApJ, 250, 478
- La Mura et al. (2007) La Mura, G., Popović, L. Č., Ciroi, S., Rafanelli, P., & Ilić, D. 2007, ApJ, 671, 104
- Lai et al. (2023) Lai, S., Wolf, C., Onken, C. A., & Bian, F. 2023, MNRAS, 521, 3682
- Lambrides et al. (2024) Lambrides, E., Garofali, K., Larson, R., et al. 2024, arXiv e-prints, arXiv:2409.13047
- Larson et al. (2023) Larson, R. L., Finkelstein, S. L., Kocevski, D. D., et al. 2023, ApJ, 953, L29
- Le Fèvre et al. (2020) Le Fèvre, O., Béthermin, M., Faisst, A., et al. 2020, A&A, 643, A1
- Lira et al. (2018) Lira, P., Kaspi, S., Netzer, H., et al. 2018, ApJ, 865, 56
- Loeb & Barkana (2001) Loeb, A. & Barkana, R. 2001, ARA&A, 39, 19
- Long et al. (2023) Long, K., Dexter, J., Cao, Y., et al. 2023, ApJ, 953, 184
- Lupi et al. (2024) Lupi, A., Trinca, A., Volonteri, M., Dotti, M., & Mazzucchelli, C. 2024, A&A, 689, A128
- Maiolino et al. (2024) Maiolino, R., Scholtz, J., Curtis-Lake, E., et al. 2024, A&A, 691, A145
- Maoz et al. (1993) Maoz, D., Netzer, H., Peterson, B. M., et al. 1993, ApJ, 404, 576
- Marziani et al. (2017) Marziani, P., Negrete, C. A., Dultzin, D., et al. 2017, Frontiers in Astronomy and Space Sciences, 4, 16
- Matsuoka et al. (2007) Matsuoka, Y., Oyabu, S., Tsuzuki, Y., & Kawara, K. 2007, ApJ, 663, 781
- Matthee et al. (2024) Matthee, J., Naidu, R. P., Brammer, G., et al. 2024, ApJ, 963, 129
- McConnell & Ma (2013) McConnell, N. J. & Ma, C.-P. 2013, ApJ, 764, 184
- Mehta et al. (2024) Mehta, D., Regan, J. A., & Prole, L. 2024, The Open Journal of Astrophysics, 7, 107
- Mizumoto et al. (2021) Mizumoto, M., Nomura, M., Done, C., Ohsuga, K., & Odaka, H. 2021, MNRAS, 503, 1442
- Murray et al. (1995) Murray, N., Chiang, J., Grossman, S. A., & Voit, G. M. 1995, ApJ, 451, 498
- Nemmen & Brotherton (2010) Nemmen, R. S. & Brotherton, M. S. 2010, MNRAS, 408, 1598
- Netzer (2019) Netzer, H. 2019, MNRAS, 488, 5185
- Netzer (2020) Netzer, H. 2020, MNRAS, 494, 1611
- Netzer & Davidson (1979) Netzer, H. & Davidson, K. 1979, MNRAS, 187, 871
- Osterbrock (1978) Osterbrock, D. E. 1978, Proceedings of the National Academy of Science, 75, 540
- Osterbrock (1989) Osterbrock, D. E. 1989, Astrophysics of gaseous nebulae and active galactic nuclei (University Science Books)
- Pancoast et al. (2014) Pancoast, A., Brewer, B. J., & Treu, T. 2014, MNRAS, 445, 3055
- Park et al. (2022) Park, D., Barth, A. J., Ho, L. C., & Laor, A. 2022, ApJS, 258, 38
- Park et al. (2013) Park, D., Woo, J.-H., Denney, K. D., & Shin, J. 2013, ApJ, 770, 87
- Prince et al. (2023) Prince, R., Zajaček, M., Panda, S., et al. 2023, A&A, 678, A189
- Ramolla et al. (2018) Ramolla, M., Haas, M., Westhues, C., et al. 2018, A&A, 620, A137
- Ramos Almeida & Ricci (2017) Ramos Almeida, C. & Ricci, C. 2017, Nature Astronomy, 1, 679
- Reines et al. (2013) Reines, A. E., Greene, J. E., & Geha, M. 2013, ApJ, 775, 116
- Richards et al. (2006) Richards, G. T., Lacy, M., Storrie-Lombardi, L. J., et al. 2006, ApJS, 166, 470
- Robertson (2022) Robertson, B. E. 2022, ARA&A, 60, 121
- Runnoe et al. (2013) Runnoe, J. C., Brotherton, M. S., Shang, Z., & DiPompeo, M. A. 2013, MNRAS, 434, 848
- Saglia et al. (2016) Saglia, R. P., Opitsch, M., Erwin, P., et al. 2016, ApJ, 818, 47
- Santos et al. (2025) Santos, D. J. D., Shimizu, T., Davies, R., et al. 2025, A&A, 696, A30
- Scholtz et al. (2025) Scholtz, J., Maiolino, R., D’Eugenio, F., et al. 2025, A&A, 697, A175
- Shen et al. (2024) Shen, Y., Grier, C. J., Horne, K., et al. 2024, ApJS, 272, 26
- Shen & Ho (2014) Shen, Y. & Ho, L. C. 2014, Nature, 513, 210
- Shimizu et al. (2019) Shimizu, T. T., Davies, R. I., Lutz, D., et al. 2019, MNRAS, 490, 5860
[…]
- Somerville & Davé (2015) Somerville, R. S. & Davé, R. 2015, ARA&A, 53, 51
- Speagle (2020) Speagle, J. S. 2020, MNRAS, 493, 3132
- Suh et al. (2025) Suh, H., Scharwächter, J., Farina, E. P., et al. 2025, Nature Astronomy, 9, 271
- Sulentic & Marziani (2015) Sulentic, J. & Marziani, P. 2015, Frontiers in Astronomy and Space Sciences, 2, 6
- Sulentic et al. (2011) Sulentic, J., Marziani, P., & Zamfir, S. 2011, Baltic Astronomy, 20, 427
- Tacconi et al. (2020) Tacconi, L. J., Genzel, R., & Sternberg, A. 2020, ARA&A, 58, 157
- Temple et al. (2023) Temple, M. J., Matthews, J. H., Hewett, P. C., et al. 2023, MNRAS, 523, 646
- Tsuzuki et al. (2006) Tsuzuki, Y., Kawara, K., Yoshii, Y., et al. 2006, ApJ, 650, 57
- Übler et al. (2024) Übler, H., D’Eugenio, F., Perna, M., et al. 2024, MNRAS, 533, 4287
- Übler et al. (2023) Übler, H., Maiolino, R., Curtis-Lake, E., et al. 2023, A&A, 677, A145
- Vayner et al. (2024) Vayner, A., Zakamska, N. L., Ishikawa, Y., et al. 2024, ApJ, 960, 126
- Véron-Cetty et al. (2004) Véron-Cetty, M. P., Joly, M., & Véron, P. 2004, A&A, 417, 515
- Vestergaard & Peterson (2006) Vestergaard, M. & Peterson, B. M. 2006, ApJ, 641, 689
- Vietri et al. (2020) Vietri, G., Mainieri, V., Kakkad, D., et al. 2020, A&A, 644, A175
- Vietri et al. (2018) Vietri, G., Piconcelli, E., Bischetti, M., et al. 2018, A&A, 617, A81
- Vollmer et al. (2018) Vollmer, B., Schartmann, M., Burtscher, L., et al. 2018, A&A, 615, A164
- Ward et al. (2024) Ward, S. R., Costa, T., Harrison, C. M., & Mainieri, V. 2024, MNRAS, 533, 1733
- Watarai (2006) Watarai, K.-y. 2006, ApJ, 648, 523
- Wolf et al. (2024) Wolf, C., Lai, S., Onken, C. A., et al. 2024, Nature Astronomy, 8, 520
- Woo et al. (2024) Woo, J.-H., Wang, S., Rakshit, S., et al. 2024, ApJ, 962, 67
- Woo et al. (2015) Woo, J.-H., Yoon, Y., Park, S., Park, D., & Kim, S. C. 2015, ApJ, 801, 38
- Wright (2006) Wright, E. L. 2006, PASP, 118, 1711
- Wylezalek et al. (2022) Wylezalek, D., Vayner, A., Rupke, D. S. N., et al. 2022, ApJ, 940, L7
- Yong et al. (2020) Yong, S. Y., Webster, R. L., King, A. L., et al. 2020, MNRAS, 491, 1320
- Yong et al. (2017) Yong, S. Y., Webster, R. L., King, A. L., et al. 2017, PASA, 34, e042
- Zubovas & King (2012) Zubovas, K. & King, A. 2012, ApJ, 745, L34
- Zubovas & Nardini (2020) Zubovas, K. & Nardini, E. 2020, MNRAS, 498, 3633
Appendix A Details of the model fit
Parameter | Short description | Priora | Value | 1 range | |
global geometry | |||||
inclination (from face-on) [deg] | U(: 0.5,1.0)b | 33.3 | -2.9,+3.0 | ||
position angle of line of nodes (E of N) [deg] | U(-180,180) | 98.5 | -3.8,+4.4 | ||
black hole mass (M⊙) | U(7.0,11.0) | 8.90 | -0.13,+0.11 | ||
fraction of all the clouds in the outflowing part | U(0.0,1.0) | 0.83 | -0.04,+0.03 | ||
geometry of disky part | |||||
angular extent above/below mid-plane [deg] | U(0,90) | 38.9 | -3.6,+5.3 | ||
geometry of outflowing part | |||||
angular extent above/below disky part [deg] | U(0,90) | 21.1 | -5.6,+6.0 | ||
scaling factor applied to at to infer | U(: 0.0,0.7) | 2.70 | -0.39,+0.47 | ||
index denoting radial dependence of | U(-1.0,0.0) | -0.74 | -0.19,+0.19 | ||
global line emission properties | |||||
shift of zero velocity wavelength, | U(-0.01,0.01) | 0.0046 | -0.0005,+0.0004 | ||
minimum radius of line emission [Schwarzschild radii] | U(: 3.0,6.0) | 3.77 | -0.17,+0.17 | ||
line emission anisotropy (-0.5 to +0.5; 0 denotes isotropic) | U(-0.5,0.5) | -0.45 | -0.03,+0.06 | ||
mid-plane opacity (0 to 1; 0 denotes opaque) | U(0.0,1.0) | 0.016 | -0.011,+0.021 | ||
H line properties | |||||
[H] | shape parameter for radial distribution of line emission | U(0.0,2.0) | 1.01 | -0.25,+0.20 | |
[H] | scale factor such that mean radius | U(0.0,1.0) | 0.46 | -0.18,+0.14 | |
[H] | flux scaling for peak of line profile | G(0.5,0.1) | 0.510 | -0.014,-+0.020 | |
H line properties | |||||
[H] | shape parameter for radial distribution of line emission | U(0.0,2.0) | 1.39 | -0.39,+0.35 | |
[H] | scale factor such that mean radius | U(0.0,1.0) | 0.50 | -0.19,+0.18 | |
[H] | flux scaling for peak of line profile | G(0.14,0.02) | 0.132 | -0.005,+0.005 |
The parameters for the best-fit model, and their 1- ranges, are given in Table 4. Here they have been grouped to make their role in the model easier to follow. The model has been fitted to the H and H lines simultaneously. As described in Kuhn et al. (2024), in practice this is done using two models in which all the geometry parameters as well as the global line emission parameters are tied to be the same in both. The only parameters that can differ between the models are those listed specifically for the H versus H line properties. These correspond to the shape parameter , the radial scaling parameter , and the peak flux scaling .
Histograms of the posterior distributions for the parameters are shown in Fig. 9, indicating that their ranges are generally well constrained by the data. As was done by Kuhn et al. (2024), we used the Python package dynesty (Speagle 2020) to fit the data with a nested sampling algorithm. Compared to the typical Markov Chain Monte Carlo algorithm, this method is powerful in terms of handling complex models with many (e.g. ) parameters and a potentially multimodal posterior distribution. We used the dynamic nested sampler (DynamicNestedSampler) which better estimates the likelihood function by re-sampling the posterior function a few more times after the “baseline run.” We used 1000 live points for the baseline run and added 500 points for each of the re-samplings until the default stop criteria was met. We adopted the random walk algorithm (rwalk) to sample the prior space and used the multi-ellipse method (multi) to create the nest boundaries. We adopted the default values for all the remaining options of dynesty. The reported fitting ends up with 2000 live points and effective samples, which is a statistically sufficient sampling of the posterior space. A corner plot of selected parameters is presented in Fig. 10. In addition to the high-probability peak, some low-probability solutions appear due to the very detailed sampling.
The resulting fits to the spectra and differential phases for the 6 baselines are shown in Fig. 8. The strongest phase signals are seen for unit telescope (UT) baselines 4-1 and 3-1, which are the longest pair at 130 m and 102 m and differ in orientation by . A weaker signal is present in baselines 3-2 and 2-1 which are at a very similar orientation to 3-1, differing by , and about half the length at 57 m and 47 m. Very little signal is seen in the 4-3 and 4-2 baselines which are oriented in a different direction. The phase signal has two peaks, the more significant one associated with the dominant outflow signature, and a weaker one related to the disky part of the BLR.